For a car loan using simple interest at a given​ rate, the amount of interest charged varies jointly with the loan amount​ (also known as the​ principal) and the time of the loan​ (in years). If a $10,000 car loan earns ​$1680 in simple interest over 6 ​years, how much interest will a ​$24,000 car loan earn over 7 ​years?
Answer by rfer(16322)

1680/6=280
280/10000=0.028=2.8% per year
24000*0.028*7=4704